OpenStudy (anonymous):

A circular pizza with a diameter of 16 inches is cut into pieces, each of which is a sector with a central angle of 40 degrees. what is the area of one piece

OpenStudy (anonymous):

Area of a sector (given in degrees) is n/360 x (pi)(radius)2 so the area of that one piece is 40/360 (pi)(8^2) = 1/9 (pi)(64) = 64pi/9

OpenStudy (anonymous):

is the n just a variable you chose?

OpenStudy (anonymous):

In that formula, n = # of degrees in the central angle

OpenStudy (anonymous):

Thank you again you're so nice for helping me

OpenStudy (anonymous):

sure.